the sum of two numbers is 29. Find the two numbers if the larger is 11 more than twice the smaller.
Answer by stanbon(75887)

the sum of two numbers is 29. Find the two numbers if the larger is 11 more than twice the smaller.
----------------
Equations:
x + y = 29
x = 2y+11
-------------
Substitute for "x" and solve for "y":
2y+11+y = 29
------
3y = 18
y = 6
----
x = 2y+1 = 13
================
Cheers,
Stan H.